Dolphins of the open ocean are classified as Type
Answer #1

Total distance travelled = D = 2 d = 2 x 90 =180 m

V = speed of sound in water = 1530 m/s

t = time for echo = distance / Speed = D/V = 180 / 1530 = 0.12 sec

B)

wavelength = Speed / frequency = 1530 / 55000 = 0.0278 m = 27.8 mm